OpenStudy (anonymous):

A meterstick standing vertically casts a shadow 1.5 meters long on the ground. At the same time, a nearby tree casts a shadow 9 meters long on the ground. About how tall is the tree? Please Help!!!

OpenStudy (anonymous):

Someone help me please?

OpenStudy (anonymous):

If a 1 meter stick casts a 1.5, it's an extra 50% of the shadow. 1 x 0.50 = 0.5 1 + 0.5 = 1.5 By that logic a 2 meter stick would cast a 3 meter shadow, because 2 x 0.50 = 1 2 + 1 = 3 So: |Y x 0.50 = Z | Y = 2Z | Y = 2Z | Y = 2 x 3 | Y = 6 < = < < < < |Z + Y = 9 | Z + 2Z = 9 | 3Z = 9 | Z = 3 | Z = 3 Hope it helped, If someone thinks it should be diferent, please correct me.

OpenStudy (anonymous):

I'm not a teacher, or a specialist, but i believe this is how it should be done... Not 100% correct if it has got to deal with percentages though.

OpenStudy (anonymous):

Thank you soooo much FabiodaSilva!! You're a lifesaver.

OpenStudy (anonymous):

If there's any part of it you did not understand, please ask, i will explain

OpenStudy (anonymous):

Y = How tall Z = How much extra for the shadow

OpenStudy (anonymous):

I understand, the response was very descriptive and easy to undertstand... I don't know how to say thanks!

OpenStudy (anonymous):

you just did :P

OpenStudy (anonymous):

Hahaha.. :)